### v1

# Strain-Enhanced Altermagnetism in Ca3Ru2O7

Abstract
Perovskite compounds exhibit enhanced octahedral distortions coupled with strong electronic correlations, providing a promising platform to explore and tune altermagnetic (AM) order. In this work, we investigate AM phases in Ca3Ru2O7, a well-known perovskite that hosts antiferromagnetism coupled to structural degrees of freedom. We demonstrate that Ca3Ru2O7 is a Kramers antiferromagnet in its ground state. However, a Néel-type magnetic configuration reveals a P-2 d-wave AM, hosting orbital-selective altermagnetism analogous to Ca2RuO4. We further explore the effects of biaxial strain on the stability between the antiferromagnetic ground state and the AM phase. Our results suggest that the AM phase becomes more stable than the AFM phase below -2% compressive strain. Additionally, strain can either enhance or suppress AM bands, with enhancements reaching up to 9% under tensile strain. To quantify the AM band splitting, we introduce the ”Altermagnetic Merit Figure” and analyze the role of electronic localization, delocalization, and octahedral distortions in AM behavior and magnetic stability changes under strain.
1 Introduction
[…]
Due to the breaking of time-reversal symmetry, the spin-orbit interaction can induce the so-called weak ferromagnetism [6] in altermagnet via a canting of the spins or other mechanisms depending on the crystal symmetries of the material [7, 8, 9, 10, 11, 12].
Given these findings, it is natural to explore whether strain could also play a key role in tuning AM order in other correlated materials. In this regard, Ruddlesden-Popper perovskites, which host diverse quantum states, could represent an interesting class of systems for investigating strain-induced AM behavior. Since many of their electronic phases are closely linked to structural distortions, these perovskites offer a promising platform to study the emergence of AM phases in relation to their inherent structural and electronic degrees of freedom [13, 14, 15, 16].
Ruddlesden-Popper compounds of the form (Ca,Sr)n+1RunO3n+1 are well-known correlated systems whose intriguing electronic properties are strongly influenced by their symmetries and octahedral distortions. These distortions, including rotations, Jahn-Teller effects, and polar distortions, can break symmetries that transform opposite spin sublattices, potentially enabling the emergence of AM states [14]. Over the past decade, these materials have attracted significant attention due to their exotic electronic phases, including Mott insulating behavior [17], metamagnetism [18], and unconventional superconductivity [19]. Revisiting their magnetic phases in the context of AM order could provide new insights into phase transitions between ferromagnetism, antiferromagnetism, and AM, offering a framework to understand the interplay between symmetry, octahedral distortions, and magnetism in correlated systems.
Among the (Ca,Sr)n+1RunO3n+1 series, we focus on Ca3Ru2O7 (CRO) to explore AM states. CRO is a fascinating compound that undergoes magnetic AFM transitions with temperature [20, 21, 22]. Over the past decade, experiments have shown that its electronic phases can be manipulated by magnetic fields [23], pressure [24], and, more recently, strain [25]. This raises the question of whether CRO exhibits altermagnetism or under what conditions its AFM properties can be tuned toward an AM state. Moreover, very few altermagnetic materials lack inversion symmetry[26, 27]. Tuning altermagnetism in Ca3Ru2O7 would introduce another noncentrosymmetric AM material, enabling the study of the interplay between non-relativistic and relativistic spin-splitting.
In this work, we explore various magnetic configurations in the CRO system to identify those that exhibit AM order. We begin by analyzing the electronic properties and then discuss the symmetry characteristics that give rise to AM phases in CRO. Furthermore, we examine the system under strain to determine the conditions under which the AM phase can be stabilized. Our results demonstrate that the altermagnetic character can be enhanced or suppressed under biaxial tensile and compressive strain. We discuss the role of octahedral distortions and the localization/delocalization effects induced by the strain in the changes to altermagnetic band splitting. Additionally, we introduce the altermagnetic quantity (AMQ) as the “figure of merit” to quantify the degree of enhancement or suppression of the band splitting with respect to the non-strained system.
2 Computational details
We perform a theoretical analysis using the density functional theory with and without spin-orbit coupling (SOC). We use the plane-wave pseudopotential method implemented in the Vienna ab-initio simulation package (VASP)[28] within the generalized gradient approximations (GGA), and it gives structural properties closer to the experimental values. The electronic valence considered are: Ru: 54 and O: 22, and Ca: 334. We use a plane-wave energy cutoff of 650 eV and set a Regular Monkhorst-Pack grid of 775 to perform the atomic relaxation and 11117 to perform the self-consistent calculation. We use a fine k-grid of 14147 within the tetrahedron method for the density of states. We perform the structural optimization of the unit cell until a force convergence threshold of at least 10-3 eV/Å per atom.
To account for the electronic correlation effects in the -orbitals of Ru atoms, we apply a Hubbard on-site Coulomb parameter using the Dudarev approximation [29] with U = 1.0 eV, a sufficiently accurate value for computing the electronic and structural properties of CRO, as reported in our previous study [30].
3 Results
3.1 Exploring AM states
CRO exhibits a non-centrosymmetric crystal structure with the space group Bb21m [21]. The associated point group is mm2, which includes a twofold rotation axis along the -axis (C2z), two perpendicular mirror planes parallel to the and planes, and a helicoidal translation along (E, C2z, , , 21). This point group has been associated with the exhibition of altermagnetic (AM) phases [31]. The magnetic moment arises predominantly from the Ru 4 electrons, in which, at low temperatures, the Ru atoms order ferromagnetically within the bilayers while aligning antiferromagnetically between them, forming an A-type antiferromagnetic structure where the spins are oriented along the -axis (see Fig. 1). Given this A-type antiferromagnetic order, the primitive cell of CRO contains spin-up sites, with a translational symmetry operation that maps spin-up into spin-down sites (along the -axis). Consequently, the ground state is not altermagnetic (see Fig. 1) [32].
To explore possible AM states and their stability, we analyzed CRO under different magnetic configurations. Specifically, we examined magnetic orderings commonly observed in AB3 perovskites [34], labeled as configurations A, B, C, and D (see Fig.2). Table 1 shows the energy differences relative to the ground-state configuration A, as well as some relevant structural, electronic, and magnetic properties for our study. Our results indicate that AM states appear in configurations B and C, where configuration B has an antiferromagnetic spin arrangement along the - and -axes, while configuration C shows this along the -axis. Although these magnetic states have not been experimentally observed in CRO, similar phases have been reported in doped Ca3Ru2O7 systems, such as Ti-doped Ca3(Ru1-xTix)2O7 (configuration B) [35], and in the sister compound Sr3Ru2O7 (configuration D) [36].
Notably, variations in magnetic spin arrangements lead to significant structural and electronic changes (further structural details are given in SM table S1). Specifically, FM and AFM in-plane arrangements result in metallic (A, C) and a narrow insular state (B, D), respectively. Next, we will focus on the AM phases of CRO, primarily in Configuration B.
E | E-State | M-Phase | () | |
A | 0 | Metallic | AFM | 1.44 |
B | 34.90 | N-Insulator | AM | 1.35 |
C | 36.46 | Metallic | AM | 1.38/1.45 |
D | 39.22 | N-Insulator | AFM | 1.35 |
In addition to the electronic differences, configuration B exhibits a lower magnetic moment than configuration A (details on Table 1).
In configuration B, the magnetic moments adopt a Néel-type ordering, where spin-up and spin-down orientations alternate within the -plane and along the -axis, forming spin-singlet wave functions.
This configuration promotes spatial delocalization, which is characteristic of bonding-like states. Since bonding states enhance wave function overlap, that is, increase the spatial proximity of electrons, configuration B is more sensitive to the Hubbard- term.
As a result, a band gap opens at relatively low U values, specifically . In contrast, configuration A remains metallic under the same conditions. Further analysis is provided in the Supplemental Material (Section SII).
These contrasting results demonstrate CRO’s highly correlated electronic nature, driven by the tight interplay between magnetic and structural degrees of freedom [30, 37, 38, 39]. Finally, we confirm that these characteristics are predominantly electronic, arising from the magnetic spin configuration, and are independent of volume changes or structural distortions.
We now focus on the electronic band structure of both systems. Figure 3 shows the total and partial density of states for configurations A and B. Configuration A exhibits metallic behavior, while configuration B displays a narrow gap at the Fermi level. The Ru ions, located in an octahedral environment formed by oxygen atoms, are subjected to a crystal field that splits the Ru 4 levels, raising the eg states and lowering the t2g manifold energies. In configuration B, the orbitals are nearly filled, leading to a small gap in the half-filled and bands. Notably, the orbitals are the most affected, as the spin rearrangement occurs within the -plane. This behavior is similar to that observed in Ca2RuO4
[40].
Figure 4 shows the non-relativistic band structure for the A and B magnetic configurations. The A configuration (ground state) exhibits metallic behavior, consistent with previously reported results that neglect spin-orbit coupling (SOC) effects [41, 30]. In contrast, configuration B displays a semimetallic character and a -dependent non-relativistic spin splitting, a hallmark of altermagnetic states (see Fig.4(b)). The shaded regions on the band structure indicate regions in -space where the altermagnetic spin splitting is maximal. The detailed BZ for these paths is shown on the right Fig. 4.
The presence of the non-relativistic spin-splitting is independent of the value of ky. According to the notation on the spin-momentum locking, this is P-2 d-wave altermagnets (where the d-wave is in the k-space) [32], since it is independent on ky this is a P-2 dxz-wave. These paths, which extend beyond the plane, have not been explored in previous studies, where ARPES experiments primarily focused on the plane [41]. In the next section, we discuss the preference for a specific direction in the AM character.
The bands located between eV exhibit significant non-relativistic spin splitting, corresponding to the - orbitals. In contrast, the bands between eV, dominated by the orbital character, showing negligible spin splitting (see, Fig.LABEL:fig:S2). Therefore, configuration B exhibits orbital-selective altermagnetism with the altermagnetism present in the - bands but not in the bands, this is similar to that observed in Ca2RuO4 [40]. When the or coordinate is inverted in reciprocal space, the non-relativistic spin splitting changes sign, while inversion of leaves it unchanged. Notably, this altermagnetic order matches that found in Ca2RuO4, despite its crystallization in a different space group [40]. The orbital-selective magnetism depends on the magnetic space group and orbital character. In conf.B, the nearly filled dxy orbitals do not contribute to magnetism, leading to orbital-selective AM. Consequently, conf. C displays different magnetic states in which all the t2g orbitals contribute to the AM character; therefore, it does not exhibit orbital selective AM, see more details in Fig.LABEL:fig:S2.
CRO, due to its non-centrosymmetric structure, is expected to exhibit pronounced SOC effects, as demonstrated in previous studies on configuration A (the experimental magnetic phase), where the spins are oriented along the -axis [41, 30]. In configuration B, the anisotropy also lies primarily along the -axis, with a minor component along the -axis, specifically . Figure LABEL:fig:S3 in the SM shows the band structure with SOC, by spin-component contributions and orbitals. Since the magnetic moment aligns along the -axis, a strong Kramers spin degeneracy is evident along this direction. However, spin polarization components are also observed along the -axis. The main contributors to the AM character are the dxy orbitals (from -0.7 to -0.5 eV) and the dyz orbitals (from -1.5 to -1.0 eV) (see SM, Fig. LABEL:fig:S2). The slight enhancement of the dxy contribution arises from the band splitting in the band structure induced by SOC, on top of the non-relativistic spin splitting. This induces both splitting and mixing of the spin character of the bands.
3.2 Symmetry analysis
To understand the role of symmetry in AM-CRO, consider applying a time-reversal operator (TR), which in practical terms consists of an exchange between spin-up and spin-down (see Fig. 5(a)-(b). Due to the distinct symmetry properties of many altermagnets, the original magnetic configuration cannot be recovered using only rotational symmetry operations. Following a time-reversal operation, a half-unit cell translation along the -axis, shifting (Fig. 5(c)), is combined with a screw symmetry operation by inverting and adding a half-unit cell translation along the -axis (Fig. 5(d)-(e)).
In summary, the operation z becomes necessary (Fig. 5(f)). These combined nonsymmorphic operations result in the -dependent splitting states [32]. Due to the symmetry operations that connect the spins, which are determined by rotations and translations along the and directions while leaving the direction unchanged, the AM character is found along these axes, as shown in Fig. 4. In the case of the C configuration, the symmetry operation ( +, + , - + ) becomes necessary to achieve AM order.
3.3 Tuning AFM to AM phases
To analyze the stability of configurations A and B, we calculated their total energies under various biaxial strain conditions (Fig. 6(a)). Configuration B shows the lowest energy within a compressive strain below . Fig. 6(b) shows the -lattice parameter for each strain. These results suggest a potential magnetic transition from AFM to AM states in CRO.
Next, we focus on investigating the effects of compressive (-) and tensile (+) strain on the AM features of CRO. Fig. 7 shows the band structure of configuration B under strains of = -2%, 0, and 2% (biaxial). It can be observed that compressive strain suppresses the AM features among the energy range [-2:-1] eV (compare Fig. 7 (a) with Fig.7(b)). Moreover, the tensile strain enhances the splitting being shifted to higher energies.
To quantify the effect of strain on the altermagnetic features of CRO, we calculate the average difference between spin-up and spin-down eigenvalues for each occupied band along the entire -path. Additionally, we define the altermagnetic quantity (AMQ) as the altermagnetic figure of merit, which integrates over the Brillouin zone. The AMQ is a numerical measure of the overall altermagnetic character of a system, providing a quantitative basis for comparing different strain conditions.
In our implementation, we read the spin-resolved eigenvalues from the EIGENVAL file generated by VASP, then compute the -dependent average splitting as the cumulative sum over occupied states along a particular set of k points (over a -path or -grid), given by:
where Nk is the numbers of bands at -point within specified energy range and Nocc is the occupied bands.
Note that the integration to obtain the AMQ can be performed either over the entire Brillouin zone or along a specific -path. We have verified that both methods yield similar solutions up to a constant multiplicative factor, which cancels out when percentage changes are considered.
We sum up all the occupied states rather than comparing only a few selected bands. This approach proves more reliable because the most strongly split bands can vary across different systems and even within the same material under strain, depending on details such as band character and crystal symmetry. Consequently, integrating over the entire set of occupied bands furnishes a more robust assessment of the altermagnetic spin-splitting.
The AMQ quantifies the spin-splitting across momentum space by integrating :
in our code, Simpson’s rule is used for numerical integration. When points are expressed in inverse angstroms (), the resulting AMQ is in units of . Furthermore, this straightforward procedure for quantifying the altermagnetic character may fail to correctly identify the spin-splitting in cases where bands cross or overlap. A potentially improved approach would incorporate a disentanglement method based on band character analysis, for instance, using Wannier functions or explicitly tracking the orbital character of each band. Such an approach would enable a more rigorous and accurate splitting assignment; however, its detailed implementation is beyond the scope of the present study and remains a subject for future work.
Since our primary goal is to evaluate strain effects on the altermagnetic state, we compare the AMQ under strain to the unstrained scenario. We define the percent variation as:
This relative measure properly captures how strain modifies the degree of altermagnetic spin-splitting throughout the Brillouin zone.
Fig. S9 illustrates the color-map along the entire -path (see the BZ in Fig.4) and by occupied bands (number labels at the -axis) corresponding to each strain value. The splitting by bands clearly varies as a function of the applied strain, with some bands becoming more or less altermagnetic. Under tensile strain (up to 4% of the maximum value explored), increases significantly compared to the unstrained system. By integrating using Eq. 2, we obtain AQM values of 0.0010, 0.0012, and 0.0013 for strains of -2%, 0%, and 2%, respectively. Fig.6(c) shows the percentual variation of AQM across different strain values given by Eq. 3. Under biaxial strain, the band splitting increases by approximately under tensile strain (at 2%) and decreases by about under compressive strain (at -2%).
The non-relativistic spin-splitting in Ca3Ru2O7 arises from the interplay between hopping parameters and the on-site energy difference between majority and minority -electrons. Strain introduces competing effects that influence these parameters: compression along one axis promotes electronic delocalization, weakening magnetism and reducing the AM band splitting, whereas tensile strain favors localization, enhancing magnetic exchange and increasing the AM band splitting.
From the energy–strain relationship shown in Fig. 6(a), we observe that configuration B (the AM phase) becomes energetically more favorable than configuration A (the AFM phase) under biaxial compressive strain (). Compressive strain decreases the in-plane Ru-O bond lengths, forcing an expansion along the -axis. This expansion modifies the balance of structural distortions so that configuration B is stabilized relative to configuration A. At the electronic level, the in-plane compression broadens the bands, weakening the magnetic exchange interactions and consequently reducing the AM band splitting. Conversely, under tensile strain, the lattice is stretched in the -plane while slightly contracting along , effectively narrowing the bandwidth and reinforcing the altermagnetic character (see, Fig.S8 SM). This effect is reflected in the rising AMQ values in Fig. 6(c) RS line (relaxed structures), demonstrating how strain can selectively enhance or suppress altermagnetic character.
To further assess the role of structural distortions, we performed an additional set of calculations where the atomic positions remained fixed at their zero-strain configuration. At the same time, only the lattice parameters were modified (fixed structure (FS) calculations). The results, depicted by the FS line in Fig. 6(c), reveal a similar overall trend: AMQ increases under tensile strain and decreases under compressive strain. However, for , the deviation between the relaxed and non-relaxed calculations becomes more pronounced, suggesting that octahedral distortions () play a secondary but non-negligible role in the strain response of the AM phase. Specifically, as seen in Fig. 6(d), increases under compressive strain for configuration B, further enhancing localization effects that help sustain the AM band splitting.
A direct comparison between relaxed structure and the non-relaxed structure allows us to disentangle the effects of strain-induced delocalization/localization from those arising purely due to octahedral distortions. In particular, the density of states (DOS) in Fig. S8 confirms that compressive strain broader states, reducing AM features. In contrast, tensile strain narrows the bandwidth, reinforcing the AM character. Moreover, the band structures shown in Fig. S10 illustrate how freezing the atomic positions primarily alters the gap structure rather than the AM spin splitting itself, reinforcing the observation that the observed strain-dependent AM behavior is primarily dictated by electronic exchange interactions, with octahedral distortions acting as a secondary fine-tuning mechanism.
Furthermore, an interesting behavior can be observed in the evolution of the distortion angle () as a function of strain (see Fig. 6(d)). In configuration A, increases monotonically from to strain. However, in configuration B, exhibits a more complex response, increasing under compressive strain () but displaying a non-monotonic behavior under tensile strain. This difference originates from the distinct in-plane magnetic configurations of the two phases. Compressive strain generally enhances electronic delocalization, which tends to weaken magnetic exchange interactions and favor a more symmetric crystal structure. In the case of conf. A, which is metallic (see, Fig. 3a)), the effect of in-plane strain gradually increases delocalization, reducing structural distortions and stabilizing a more symmetric octahedral environment. Conversely, electronic states are more localized in configuration B (see, Fig. 3b)), where the in-plane magnetic alignment is AFM. As a result, under compression, the system seeks to accommodate strain by increasing octahedral distortions (), effectively enhancing localization to maintain magnetic interactions, resulting in a more distorted structure than configuration A. Under tensile strain, however, competing effects emerge: while in-plane expansion reduces the need for further distortions, the slight contraction along the -axis counterbalances this effect, leading to either an increase or decrease in , depending on the precise strain conditions.
The strain-dependent behavior of is closely tied to the observed changes in the energy profile in Fig. 6(a), where configuration B becomes energetically favorable at compressive strain values above . The stabilization of the AM phase coincides with an increase in , suggesting that octahedral distortions play a crucial role in the strain-driven AFM-to-AM transition.
4 Final Remarks
In this work, we explore the emergence of altermagnetic states in the correlated material Ca3Ru2O7. While its ground state lacks AM features, alternative magnetic configurations enable their realization. AM arises in spin arrangements connected through rotational and translational symmetries. In particular, configuration B, which alternates spin orientation both in-plane and between layers, exhibits spin connectivity via rotations along the -axis and translations in the - plane. This configuration hosts a P-2 -wave altermagnetic state, which is orbital-selective, with AM primarily arising from the - orbitals. When the spin-orbit coupling is included, with the Néel vector oriented along the -axis, time-reversal symmetry breaking leads to weak ferromagnetism along the -axis due to the staggered Dzyaloshinskii-Moriya interaction [7, 15]. While configuration B has not been experimentally observed in pristine Ca3Ru2O7, it has been reported in doped systems, such as Ca3(TixRu1-x)2O7, suggesting that this system could host an AM ground state under suitable conditions. Additionally, our results indicate that strain plays a crucial role in tuning AM properties. Tensile strain along the -axis enhances AM band splitting by approximately (for strain), whereas compressive strain suppresses it by . Moreover, we predict a strain-driven phase transition from an AFM to an AM state (configuration A to B) under compressive strain exceeding . While previous experiments have manipulated magnetic anisotropy using uniaxial strain (), the impact of biaxial strain remains largely unexplored.
Strain-tuned AM phases could significantly expand the known quantum phenomena in the Can+1RunO3n+1 family, which already exhibit strain-driven transitions, such as Mott phase suppression in Ca2RuO4 [17], the emergence of a Kondo effect, and a metal-to-semiconductor transition in CaRuO3 ( strain) [42]. Our findings provide a fundamental step toward uncovering possible AM states in the Ca3Ru2O7 family and pave the way for future theoretical and experimental investigations of biaxial strain as a tool for engineering AM phases in correlated oxides.
Acknowledgments
A.L. thanks the Technische Universität Dresden Professor Fellowship Program. C. A. acknowledges the useful discussion with G. Cuono and A. Fakhredine about altermagnetism in Ca2RuO4. C. A. is supported by the Foundation for Polish Science project “MagTop” no. FENG.02.01-IP.05-0028/23 co-financed by the European Union from the funds of Priority 2 of the European Funds for a Smart Economy Program 2021–2027 (FENG). J.W.G. acknowledges financial support from Chilean ANID-FONDECY grants N. 1220700 and 1221301. T.B. would like to thank the German Science Foundation (DFG) for supporting this work via the German-Israeli Project Cooperation with the Grant No. HE 3543/42-1.

### v2

@@
-Strain-Enhanced Altermagnetism in Ca3Ru2O7
-Abstract
-Perovskite compounds exhibit enhanced octahedral distortions coupled with strong electronic correlations, providing a promising platform to explore and tune altermagnetic (AM) order.
-In this work, we investigate AM phases in Ca3Ru2O7, a well-known perovskite that hosts antiferromagnetism coupled to structural degrees of freedom.
-We demonstrate that Ca3Ru2O7 is a Kramers antiferromagnet in its ground state.
-However, a Néel-type magnetic configuration reveals a P-2 d-wave AM, hosting orbital-selective altermagnetism analogous to Ca2RuO4.
-We further explore the effects of biaxial strain on the stability between the antiferromagnetic ground state and the AM phase.
-Our results suggest that the AM phase becomes more stable than the AFM phase below -2% compressive strain.
-Additionally, strain can either enhance or suppress AM bands, with enhancements reaching up to 9% under tensile strain.
-To quantify the AM band splitting, we introduce the ”Altermagnetic Merit Figure” and analyze the role of electronic localization, delocalization, and octahedral distortions in AM behavior and magnetic stability changes under strain.
-1 Introduction
-[…]
-Due to the breaking of time-reversal symmetry, the spin-orbit interaction can induce the so-called weak ferromagnetism [6] in altermagnet via a canting of the spins or other mechanisms depending on the crystal symmetries of the material [7, 8, 9, 10, 11, 12].
-Given these findings, it is natural to explore whether strain could also play a key role in tuning AM order in other correlated materials.
-In this regard, Ruddlesden-Popper perovskites, which host diverse quantum states, could represent an interesting class of systems for investigating strain-induced AM behavior.
-Since many of their electronic phases are closely linked to structural distortions, these perovskites offer a promising platform to study the emergence of AM phases in relation to their inherent structural and electronic degrees of freedom [13, 14, 15, 16].
-Ruddlesden-Popper compounds of the form (Ca,Sr)n+1RunO3n+1 are well-known correlated systems whose intriguing electronic properties are strongly influenced by their symmetries and octahedral distortions.
-These distortions, including rotations, Jahn-Teller effects, and polar distortions, can break symmetries that transform opposite spin sublattices, potentially enabling the emergence of AM states [14].
-Over the past decade, these materials have attracted significant attention due to their exotic electronic phases, including Mott insulating behavior [17], metamagnetism [18], and unconventional superconductivity [19].
-Revisiting their magnetic phases in the context of AM order could provide new insights into phase transitions between ferromagnetism, antiferromagnetism, and AM, offering a framework to understand the interplay between symmetry, octahedral distortions, and magnetism in correlated systems.
-Among the (Ca,Sr)n+1RunO3n+1 series, we focus on Ca3Ru2O7 (CRO) to explore AM states.
-CRO is a fascinating compound that undergoes magnetic AFM transitions with temperature [20, 21, 22].
-Over the past decade, experiments have shown that its electronic phases can be manipulated by magnetic fields [23], pressure [24], and, more recently, strain [25].
-This raises the question of whether CRO exhibits altermagnetism or under what conditions its AFM properties can be tuned toward an AM state.
-Moreover, very few altermagnetic materials lack inversion symmetry[26, 27].
-Tuning altermagnetism in Ca3Ru2O7 would introduce another noncentrosymmetric AM material, enabling the study of the interplay between non-relativistic and relativistic spin-splitting.
-In this work, we explore various magnetic configurations in the CRO system to identify those that exhibit AM order.
-We begin by analyzing the electronic properties and then discuss the symmetry characteristics that give rise to AM phases in CRO.
-Furthermore, we examine the system under strain to determine the conditions under which the AM phase can be stabilized.
-Our results demonstrate that the altermagnetic character can be enhanced or suppressed under biaxial tensile and compressive strain.
-We discuss the role of octahedral distortions and the localization/delocalization effects induced by the strain in the changes to altermagnetic band splitting.
-Additionally, we introduce the altermagnetic quantity (AMQ) as the “figure of merit” to quantify the degree of enhancement or suppression of the band splitting with respect to the non-strained system.
-2 Computational details
-We perform a theoretical analysis using the density functional theory with and without spin-orbit coupling (SOC).
-We use the plane-wave pseudopotential method implemented in the Vienna ab-initio simulation package (VASP)[28] within the generalized gradient approximations (GGA), and it gives structural properties closer to the experimental values.
-The electronic valence considered are: Ru: 54 and O: 22, and Ca: 334.
-We use a plane-wave energy cutoff of 650 eV and set a Regular Monkhorst-Pack grid of 775 to perform the atomic relaxation and 11117 to perform the self-consistent calculation.
-We use a fine k-grid of 14147 within the tetrahedron method for the density of states.
-We perform the structural optimization of the unit cell until a force convergence threshold of at least 10-3 eV/Å per atom.
-To account for the electronic correlation effects in the -orbitals of Ru atoms, we apply a Hubbard on-site Coulomb parameter using the Dudarev approximation [29] with U = 1.0 eV, a sufficiently accurate value for computing the electronic and structural properties of CRO, as reported in our previous study [30].
-3 Results
-3.1 Exploring AM states
-CRO exhibits a non-centrosymmetric crystal structure with the space group Bb21m [21].
-The associated point group is mm2, which includes a twofold rotation axis along the -axis (C2z), two perpendicular mirror planes parallel to the and planes, and a helicoidal translation along (E, C2z, , , 21).
-This point group has been associated with the exhibition of altermagnetic (AM) phases [31].
+Hybrid d/p-wave altermagnetism in Ca3Ru2O7 and strain-controlled spin splitting
+Hybrid d/p-wave altermagnetism in Ca3Ru2O7 and strain-controlled spin splitting
+2Dresden University of Technology, Institute for Solid State and Materials Physics, 01062 Dresden, Germany
+3International Research Centre Magtop, Institute of Physics, Polish Academy of Sciences, Aleja Lotników 32/46, 02668 Warsaw, Poland.
+4Chair of Theoretical Chemistry, Technische Universität Dresden, Bergstrasse 66, 01069 Dresden, Germany.
+5Departamento de Física, Universidad de Antofagasta, Av. Angamos 601, Casilla 170, Antofagasta, Chile.
+The interplay of strong electronic correlations, sizable octahedral distortions, and pronounced spin–orbit coupling (SOC) makes perovskite oxides promising candidates for realizing altermagnetic phases.
+We study altermagnetic phases in Ca3Ru2O7, a non-centrosymmetric layered perovskite whose ground state is a Kramers-degenerate antiferromagnet.
+We show that an alternative Néel-type spin arrangement hosts a P-2 -wave altermagnetic state with orbital selectivity similar to Ca2RuO4.
+Including SOC generates a symmetry-allowed -wave component and yields a hybrid /-wave altermagnetic order.
+We further demonstrate that biaxial strain tunes both magnetic stability and band splitting: compressive strain beyond 2% favors the altermagnetic phase over the antiferromagnetic ground state, while tensile strain increases altermagnetic splittings by up to 9%.
+To quantify these trends, we define an altermagnetic figure of merit and trace its strain dependence to changes in electronic localization and octahedral geometry in this polar metal.
+Introduction
+11footnotetext: * Corresponding author: andrea.leon@uchile.cl22footnotetext: † jhon.gonzalez@uantof.clAltermagnets are a significant new magnetic material class, advancing fundamental research and spintronic applications [1, 2].
+At the heart of altermagnetism (AM) lies crystal symmetry, which dictates the unique spin-splitting behavior.
+This inherent link makes strain a powerful tool for manipulating and controlling AM order in correlated materials, a frontier of active research [3, 4].
+The potential of strain engineering has been recently demonstrated by the induced phase transition from an antiferromagnetic to an altermagnetic state in ReO2 [4] and the discovery of an elasto-Hall conductivity in d-wave altermagnets [5].
+The physics is further enriched by spin-orbit coupling (SOC), which, by breaking time-reversal symmetry, can induce phenomena such as weak ferromagnetism through spin canting or other mechanisms dependent on crystal symmetry [6, 7, 8, 9, 10, 11, 12].
+Collectively, these findings motivate the exploration of strain-induced AM order in new classes of correlated materials.
+In this context, Ruddlesden-Popper perovskites, renowned for hosting diverse quantum phases intrinsically coupled to lattice distortions, offer a compelling platform to investigate the interplay between strain, symmetry, and magnetism [13, 14, 15, 16, 17].
+Ruddlesden-Popper compounds of the form (Ca, Sr)n+1RunO3n+1 exhibit strong electronic correlations and display phases that depend critically on symmetry and octahedral distortions.
+Rotations, Jahn-teller distortions, and polar displacements break the operations that interchange opposite spin sublattices and enable the emergence of altermagnetic states [16].
+Over the past decade, researchers have observed Mott insulating behavior [18], metamagnetism [19], and unconventional superconductivity [20] in these systems.
+Revisiting their magnetic phases within the altermagnetic framework offers new insights into transitions among ferromagnetism, antiferromagnetism, and altermagnetism by linking symmetry, distortion, and magnetic order.
+Among the series, Ca3Ru2O7 undergoes antiferromagnetic transitions as a function of temperature [21, 22, 23] and exhibits pronounced tunability under magnetic field [24], pressure [25] and strain [26].
+These observations raise the question of whether Ca3Ru2O7 (hereafter CRO) hosts an altermagnetic phase and what conditions might induce a transition from its native antiferromagnetic order.
+We apply this framework to Ca3Ru2O7, a non-centrosymmetric compound with strong spin-orbit coupling (SOC) that enables relativistic spin splitting and spin-momentum locking.
+SOC governs the magnetic anisotropy in this system and drives spin reorientation transitions, as previously reported [27].
+Investigating altermagnetism in Ca3Ru2O7 offers a unique opportunity to reveal spin textures that emerge from the interplay of non-relativistic and relativistic effects.
+Most known altermagnets preserve inversion symmetry; finding altermagnetic order in a non-centrosymmetric metal would open new directions in the study of symmetry-protected spin transport phenomena [28, 29].
+In this work, we identify magnetic configurations in Ca3Ru2O7 that host altermagnetic order and analyze their symmetry and electronic structure under non-relativistic and relativistic spin splitting.
+In the absence of SOC, the system exhibits a -wave altermagnetic state, while SOC introduces additional asymmetries and generates a hybrid - and -wave character.
+We demonstrate that biaxial strain tunes the altermagnetic response by enhancing or suppressing spin splitting depending on the strain direction and link this manipulation to changes in octahedral distortions and orbital localization.
+To quantify these effects, we define the altermagnetic quantity (AMQ) as a figure of merit that captures changes in band splitting and enables direct comparison across altermagnetic systems across different conditions.
+Results
+Exploring AM states
+Ca3Ru2O7 crystallizes in the non-centrosymmetric space group Bb21m [22], with associated point group mm2, which includes a twofold rotation axis along and a vertical mirror plane.
+The full space group also contains a 21 screw axis along , which plays a role in the nonsymmorphic symmetries relevant to the altermagnetic configuration.
+[30].
 The magnetic moment arises predominantly from the Ru 4 electrons, in which, at low temperatures, the Ru atoms order ferromagnetically within the bilayers while aligning antiferromagnetically between them, forming an A-type antiferromagnetic structure where the spins are oriented along the -axis (see Fig.
@@
 1) [31].
-To explore possible AM states and their stability, we analyzed CRO under different magnetic configurations.
-Specifically, we examined magnetic orderings commonly observed in AB3 perovskites [34], labeled as configurations A, B, C, and D (see Fig.2).
+We analyzed CRO under different magnetic configurations to explore possible AM states and their stability.
+Specifically, we examined magnetic orderings commonly observed in AB3 perovskites [33], labeled as configurations A, B, C, and D (see Fig.
+2).
 Table 1 shows the energy differences relative to the ground-state configuration A, as well as some relevant structural, electronic, and magnetic properties for our study.
@@
 Although these magnetic states have not been experimentally observed in CRO, similar phases have been reported in doped Ca3Ru2O7 systems, such as Ti-doped Ca3(Ru1-xTix)2O7 (configuration B) [34], and in the sister compound Sr3Ru2O7 (configuration D) [35].
-Notably, variations in magnetic spin arrangements lead to significant structural and electronic changes (further structural details are given in SM table S1).
+Notably, variations in magnetic spin arrangements lead to significant structural and electronic changes (further structural details are given in supplementary information file (SI) table S1).
 Specifically, FM and AFM in-plane arrangements result in metallic (A, C) and a narrow insular state (B, D), respectively.
 Next, we will focus on the AM phases of CRO, primarily in Configuration B.
-E | E-State | M-Phase | () | |
-A | 0 | Metallic | AFM | 1.44 |
-B | 34.90 | N-Insulator | AM | 1.35 |
-C | 36.46 | Metallic | AM | 1.38/1.45 |
-D | 39.22 | N-Insulator | AFM | 1.35 |
-In addition to the electronic differences, configuration B exhibits a lower magnetic moment than configuration A (details on Table 1).
-In configuration B, the magnetic moments adopt a Néel-type ordering, where spin-up and spin-down orientations alternate within the -plane and along the -axis, forming spin-singlet wave functions.
-This configuration promotes spatial delocalization, which is characteristic of bonding-like states.
-Since bonding states enhance wave function overlap, that is, increase the spatial proximity of electrons, configuration B is more sensitive to the Hubbard- term.
+In addition to the electronic differences, configuration B exhibits a lower magnetic moment than configuration A (details on Table S1).
+In configuration B, the magnetic moments form a Néel-type collinear ordering with alternating spin orientations.
+This promotes spin-compensated orbital overlap and wave function delocalization, which are characteristic of bonding-like states.
+As a result, configuration B is more sensitive to the Hubbard- term due to the enhanced electronic proximity.
 As a result, a band gap opens at relatively low U values, specifically .
 In contrast, configuration A remains metallic under the same conditions.
-Further analysis is provided in the Supplemental Material (Section SII).
+The Supplemental Material (Section SII) provides further analysis.
 These contrasting results demonstrate CRO’s highly correlated electronic nature, driven by the tight interplay between magnetic and structural degrees of freedom [36, 37, 38, 39].
 Finally, we confirm that these characteristics are predominantly electronic, arising from the magnetic spin configuration, and are independent of volume changes or structural distortions.
-We now focus on the electronic band structure of both systems.
-Figure 3 shows the total and partial density of states for configurations A and B.
-Configuration A exhibits metallic behavior, while configuration B displays a narrow gap at the Fermi level.
-The Ru ions, located in an octahedral environment formed by oxygen atoms, are subjected to a crystal field that splits the Ru 4 levels, raising the eg states and lowering the t2g manifold energies.
-In configuration B, the orbitals are nearly filled, leading to a small gap in the half-filled and bands.
-Notably, the orbitals are the most affected, as the spin rearrangement occurs within the -plane.
-This behavior is similar to that observed in Ca2RuO4
-[40].
+We now examine the electronic structure of both configurations.
+Figure 3 shows the total and orbital-resolved density of states.
+While configuration A is metallic, configuration B exhibits a narrow gap at the Fermi level.
+In an ideal octahedral environment, the Ru , , and orbitals are degenerate.
+Orthorhombic distortions lift this degeneracy slightly, but it is the Néel-type magnetic order in the -plane (conf.
+B) that drives orbital differentiation by enhancing -type hybridization between Ru- and O- orbitals.
+This leads to a strong bonding–antibonding splitting, pushing the bonding states to lower energy and making them nearly filled.
+The resulting spectral redistribution lowers the chemical potential relative to the less-affected, half-filled and orbitals.
+The gap thus reflects an orbital-selective mechanism driven by the interplay between magnetism and crystal-field effects, as proposed for Ca2RuO4 [40]
 Figure 4 shows the non-relativistic band structure for the A and B magnetic configurations.
-The A configuration (ground state) exhibits metallic behavior, consistent with previously reported results that neglect spin-orbit coupling (SOC) effects [41, 30].
-In contrast, configuration B displays a semimetallic character and a -dependent non-relativistic spin splitting, a hallmark of altermagnetic states (see Fig.4(b)).
+The A configuration (ground state) exhibits metallic behavior, consistent with previously reported results that neglect SOC effects [27, 36].
+In contrast, configuration B displays a narrow gap and a -dependent non-relativistic spin splitting, a hallmark of altermagnetic states (see Fig.
+4(b)).
 The shaded regions on the band structure indicate regions in -space where the altermagnetic spin splitting is maximal.
-The detailed BZ for these paths is shown on the right Fig.
-4.
+Figure 5 shows the detailed Brillouin zone (BZ) corresponding to these paths.
 The presence of the non-relativistic spin-splitting is independent of the value of ky.
-According to the notation on the spin-momentum locking, this is P-2 d-wave altermagnets (where the d-wave is in the k-space) [32], since it is independent on ky this is a P-2 dxz-wave.
-These paths, which extend beyond the plane, have not been explored in previous studies, where ARPES experiments primarily focused on the plane [41].
-In the next section, we discuss the preference for a specific direction in the AM character.
+The observed spin splitting exhibits what is known as a d-wave symmetry in momentum space.
+This terminology refers to how the splitting changes sign across the Brillouin zone, in analogy to the lobes of a d-orbital.
+Specifically, the splitting changes sign upon inversion of the kx or kz coordinate, but remains unchanged when inverting ky.
+This independence from the ky component allows for a more precise classification as a dxz-wave character.
+According to the formal symmetry classification of altermagnets [31], this behavior corresponds to a P-2 dxz-wave state.
+These paths, which extend beyond the xy plane, have not been explored in previous studies, where ARPES experiments primarily focused on the xy plane [27, 41].
+The following section discusses the preference for a specific direction in the AM character.
 The bands located between eV exhibit significant non-relativistic spin splitting, corresponding to the - orbitals.
-In contrast, the bands between eV, dominated by the orbital character, showing negligible spin splitting (see, Fig.LABEL:fig:S2).
+In contrast, the bands between eV, dominated by the orbital character, show negligible spin splitting (see, Fig.
+S3).
 Therefore, configuration B exhibits orbital-selective altermagnetism with the altermagnetism present in the - bands but not in the bands, this is similar to that observed in Ca2RuO4 [40].
@@
 Consequently, conf.
-C displays different magnetic states in which all the t2g orbitals contribute to the AM character; therefore, it does not exhibit orbital selective AM, see more details in Fig.LABEL:fig:S2.
-CRO, due to its non-centrosymmetric structure, is expected to exhibit pronounced SOC effects, as demonstrated in previous studies on configuration A (the experimental magnetic phase), where the spins are oriented along the -axis [41, 30].
-In configuration B, the anisotropy also lies primarily along the -axis, with a minor component along the -axis, specifically .
-Figure LABEL:fig:S3 in the SM shows the band structure with SOC, by spin-component contributions and orbitals.
-Since the magnetic moment aligns along the -axis, a strong Kramers spin degeneracy is evident along this direction.
-However, spin polarization components are also observed along the -axis.
-The main contributors to the AM character are the dxy orbitals (from -0.7 to -0.5 eV) and the dyz orbitals (from -1.5 to -1.0 eV) (see SM, Fig.
-LABEL:fig:S2).
-The slight enhancement of the dxy contribution arises from the band splitting in the band structure induced by SOC, on top of the non-relativistic spin splitting.
-This induces both splitting and mixing of the spin character of the bands.
-3.2 Symmetry analysis
+C displays different magnetic states in which all the t2g orbitals contribute to the AM character; therefore, it does not exhibit orbital selective AM, see more details in Fig.
+S3.
+The lack of inversion symmetry in Ca3Ru2O7 enables sizable SOC effects, which play a central role in its magnetic properties [27].
+In configuration B, as in configuration A, the spins align primarily along the -axis, with a slight canting toward -axis ().
+SOC enriches the spin textures and activates symmetry-breaking mechanisms that lift band degeneracies, thereby enabling access to AM states that remain hidden in the collinear case.
+Our spin-projected band structure analysis reveals how SOC qualitatively transforms the nature of altermagnetism in this system.
+The non-relativistic ground state exhibits a clear -wave character, dominated by -like states along the high-symmetry paths shown in Fig.
+4.
+Upon including SOC, a complex landscape emerges.
+The original -wave features remain most evident in the projection (Figs.
+6(a)-(b)), consistent with the magnetic anisotropy along the -axis.
+In stark contrast, the projection unveils a completely new altermagnetic signature between -0.5 and -0.75 eV, emerging from bands that were purely antiferromagnetic in the collinear case (Figs.
+6(a),(c)).
+The influence of SOC is particularly evident along the direction.
+Here, SOC lifts the band degeneracy in the channel, creating a spin splitting where the non-relativistic bands were degenerate (Figs.
+6(e)-(f)) (similar results are found for along ).
+This purely relativistic splitting exhibits a characteristic antisymmetric behavior for momentum inversion (), which is the defining hallmark of a -wave altermagnetic component.
+Therefore, these findings demonstrate the emergence of a novel hybrid -wave altermagnetic order.
+The AM+SOC phase in CRO is best described as a state where the non-relativistic -wave component coexists with a purely SOC-driven -wave character.
+This hybridization arises from the interplay between the magnetic structure and the crystal’s nonsymmorphic symmetries.
+While these results highlight a recently emerging regime of altermagnetism in the presence of SOC [42, 43], a formal classification of this hybrid order, for instance, through advanced group-theoretical analysis, remains a substantial theoretical challenge.
+we describe the emergent AM+SOC features through their direct, symmetry-resolved signatures in the spin-projected electronic structure.
+Future studies will address a more systematic classification of AM+SOC spin textures.
+Symmetry analysis
 To understand the role of symmetry in AM-CRO, consider applying a time-reversal operator (TR), which in practical terms consists of an exchange between spin-up and spin-down (see Fig.
-5(a)-(b).
+7(a)-(b).
 Due to the distinct symmetry properties of many altermagnets, the original magnetic configuration cannot be recovered using only rotational symmetry operations.
 Following a time-reversal operation, a half-unit cell translation along the -axis, shifting (Fig.
-5(c)), is combined with a screw symmetry operation by inverting and adding a half-unit cell translation along the -axis (Fig.
-5(d)-(e)).
+7(c)), is combined with a screw symmetry operation by inverting and adding a half-unit cell translation along the -axis (Fig.
+7(d)-(e)).
 In summary, the operation z becomes necessary (Fig.
-5(f)).
+7(f)).
 These combined nonsymmorphic operations result in the -dependent splitting states [31].
@@
 In the case of the C configuration, the symmetry operation ( +, + , - + ) becomes necessary to achieve AM order.
-3.3 Tuning AFM to AM phases
+This symmetry operation is specific to the orthorhombic space group Bb21m.
+Since the crystallographic axes are not symmetry-equivalent, this operation does not remain valid under the replacement
+Tuning AFM to AM phases
 To analyze the stability of configurations A and B, we calculated their total energies under various biaxial strain conditions (Fig.
-6(a)).
+8(a)).
 Configuration B shows the lowest energy within a compressive strain below .
 Fig.
-6(b) shows the -lattice parameter for each strain.
+8(b) shows the -lattice parameter for each strain.
 These results suggest a potential magnetic transition from AFM to AM states in CRO.
-Next, we focus on investigating the effects of compressive (-) and tensile (+) strain on the AM features of CRO.
+Next, we investigate the effects of compressive (-) and tensile (+) strain on the AM features of CRO.
 Fig.
@@
 It can be observed that compressive strain suppresses the AM features among the energy range [-2:-1] eV (compare Fig.
-7 (a) with Fig.7(b)).
-Moreover, the tensile strain enhances the splitting being shifted to higher energies.
+9 (a) with Fig.
+9(b)).
+Moreover, the tensile strain enhances the splitting and shifts it to higher energies.
 To quantify the effect of strain on the altermagnetic features of CRO, we calculate the average difference between spin-up and spin-down eigenvalues for each occupied band along the entire -path.
@@
 Note that the integration to obtain the AMQ can be performed either over the entire Brillouin zone or along a specific -path.
-We have verified that both methods yield similar solutions up to a constant multiplicative factor, which cancels out when percentage changes are considered.
+We verified that both methods yield similar solutions; the only difference is a constant multiplicative factor, which cancels out when computing percentage changes.
 We sum up all the occupied states rather than comparing only a few selected bands.
 This approach proves more reliable because the most strongly split bands can vary across different systems and even within the same material under strain, depending on details such as band character and crystal symmetry.
-Consequently, integrating over the entire set of occupied bands furnishes a more robust assessment of the altermagnetic spin-splitting.
+Consequently, integrating over the entire set of occupied bands offers a more robust assessment of the altermagnetic spin-splitting.
 The AMQ quantifies the spin-splitting across momentum space by integrating :
@@
 When points are expressed in inverse angstroms (), the resulting AMQ is in units of .
-Furthermore, this straightforward procedure for quantifying the altermagnetic character may fail to correctly identify the spin-splitting in cases where bands cross or overlap.
-A potentially improved approach would incorporate a disentanglement method based on band character analysis, for instance, using Wannier functions or explicitly tracking the orbital character of each band.
-Such an approach would enable a more rigorous and accurate splitting assignment; however, its detailed implementation is beyond the scope of the present study and remains a subject for future work.
+Although the AMQ is not a formal symmetry-derived order parameter like magnetic multipoles [44], it provides a symmetry-consistent and practical metric to quantify momentum-resolved spin splitting in systems with zero net magnetization.
+While this approach is robust, it may fail near band crossings or overlapping bands.
+A more rigorous treatment, such as disentanglement via Wannier-based orbital tracking, could refine the analysis but lies beyond the scope of this study.
+Nonetheless, the observed vanishing of spin splitting at high-symmetry points and the overall internal consistency strongly support the identification of altermagnetic behavior.
 Since our primary goal is to evaluate strain effects on the altermagnetic state, we compare the AMQ under strain to the unstrained scenario. We define the percent variation as:
@@
 Fig.
-S9 illustrates the color-map along the entire -path (see the BZ in Fig.4) and by occupied bands (number labels at the -axis) corresponding to each strain value.
-The splitting by bands clearly varies as a function of the applied strain, with some bands becoming more or less altermagnetic.
-Under tensile strain (up to 4% of the maximum value explored), increases significantly compared to the unstrained system.
-By integrating using Eq.
-2, we obtain AQM values of 0.0010, 0.0012, and 0.0013 for strains of -2%, 0%, and 2%, respectively.
-Fig.6(c) shows the percentual variation of AQM across different strain values given by Eq.
-3.
-Under biaxial strain, the band splitting increases by approximately under tensile strain (at 2%) and decreases by about under compressive strain (at -2%).
-The non-relativistic spin-splitting in Ca3Ru2O7 arises from the interplay between hopping parameters and the on-site energy difference between majority and minority -electrons.
+S6 shows the evolution of the spin splitting across the full -path and occupied bands under different strain conditions (see Brillouin zone in Fig.
+4).
+The magnitude of band-dependent altermagnetic splitting varies with strain, increasing notably under tensile strain.
+The AQM response varies by approximately under tensile strain (2%) and under compressive strain (-2%), as shown in Fig.
+8(c).
+All computational details and equations are provided in the Supplementary Information (Fig.
+S7).
+Discussion
+The non-relativistic spin-splitting in CRO arises from the interplay between hopping parameters and the opposite on-site spin polarizations in the two magnetic sublattices.
 Strain introduces competing effects that influence these parameters: compression along one axis promotes electronic delocalization, weakening magnetism and reducing the AM band splitting, whereas tensile strain favors localization, enhancing magnetic exchange and increasing the AM band splitting.
-From the energy–strain relationship shown in Fig.
-6(a), we observe that configuration B (the AM phase) becomes energetically more favorable than configuration A (the AFM phase) under biaxial compressive strain ().
+From the energy-strain relationship shown in Fig.
+8(a), we observe that configuration B (the AM phase) becomes energetically more favorable than configuration A (the AFM phase) under biaxial compressive strain ().
 Compressive strain decreases the in-plane Ru-O bond lengths, forcing an expansion along the -axis.
-This expansion modifies the balance of structural distortions so that configuration B is stabilized relative to configuration A.
+This expansion modifies the balance of structural distortions and stabilizes configuration B over configuration A.
 At the electronic level, the in-plane compression broadens the bands, weakening the magnetic exchange interactions and consequently reducing the AM band splitting.
-Conversely, under tensile strain, the lattice is stretched in the -plane while slightly contracting along , effectively narrowing the bandwidth and reinforcing the altermagnetic character (see, Fig.S8 SM).
+Conversely, under tensile strain, the lattice is stretched in the -plane while slightly contracting along , effectively narrowing the bandwidth and reinforcing the altermagnetic character (Fig.
+S5).
 This effect is reflected in the rising AMQ values in Fig.
-6(c) RS line (relaxed structures), demonstrating how strain can selectively enhance or suppress altermagnetic character.
+8(c) RS line (relaxed structures), demonstrating how strain can selectively enhance or suppress altermagnetic character.
 To further assess the role of structural distortions, we performed an additional set of calculations where the atomic positions remained fixed at their zero-strain configuration.
@@
 The results, depicted by the FS line in Fig.
-6(c), reveal a similar overall trend: AMQ increases under tensile strain and decreases under compressive strain.
+8(c), reveal a similar overall trend: AMQ increases under tensile strain and decreases under compressive strain.
 However, for , the deviation between the relaxed and non-relaxed calculations becomes more pronounced, suggesting that octahedral distortions () play a secondary but non-negligible role in the strain response of the AM phase.
 Specifically, as seen in Fig.
-6(d), increases under compressive strain for configuration B, further enhancing localization effects that help sustain the AM band splitting.
-A direct comparison between relaxed structure and the non-relaxed structure allows us to disentangle the effects of strain-induced delocalization/localization from those arising purely due to octahedral distortions.
-In particular, the density of states (DOS) in Fig.
-S8 confirms that compressive strain broader states, reducing AM features.
+8(d), increases under compressive strain for configuration B, further enhancing localization effects that help sustain the AM band splitting.
+A direct comparison between the relaxed structures and the non-relaxed structures allows us to disentangle the effects of strain-induced delocalization/localization from those arising purely from octahedral distortions.
+In particular, the density of states in Fig.
+S5 confirms that compressive strain broader states, reducing AM features.
 In contrast, tensile strain narrows the bandwidth, reinforcing the AM character.
 Moreover, the band structures shown in Fig.
-S10 illustrate how freezing the atomic positions primarily alters the gap structure rather than the AM spin splitting itself, reinforcing the observation that the observed strain-dependent AM behavior is primarily dictated by electronic exchange interactions, with octahedral distortions acting as a secondary fine-tuning mechanism.
-Furthermore, an interesting behavior can be observed in the evolution of the distortion angle () as a function of strain (see Fig.
-6(d)).
-In configuration A, increases monotonically from to strain.
-However, in configuration B, exhibits a more complex response, increasing under compressive strain () but displaying a non-monotonic behavior under tensile strain.
-This difference originates from the distinct in-plane magnetic configurations of the two phases.
-Compressive strain generally enhances electronic delocalization, which tends to weaken magnetic exchange interactions and favor a more symmetric crystal structure.
-In the case of conf.
-A, which is metallic (see, Fig.
-3a)), the effect of in-plane strain gradually increases delocalization, reducing structural distortions and stabilizing a more symmetric octahedral environment.
-Conversely, electronic states are more localized in configuration B (see, Fig.
-3b)), where the in-plane magnetic alignment is AFM.
-As a result, under compression, the system seeks to accommodate strain by increasing octahedral distortions (), effectively enhancing localization to maintain magnetic interactions, resulting in a more distorted structure than configuration A.
-Under tensile strain, however, competing effects emerge: while in-plane expansion reduces the need for further distortions, the slight contraction along the -axis counterbalances this effect, leading to either an increase or decrease in , depending on the precise strain conditions.
-The strain-dependent behavior of is closely tied to the observed changes in the energy profile in Fig.
-6(a), where configuration B becomes energetically favorable at compressive strain values above .
-The stabilization of the AM phase coincides with an increase in , suggesting that octahedral distortions play a crucial role in the strain-driven AFM-to-AM transition.
-4 Final Remarks
-In this work, we explore the emergence of altermagnetic states in the correlated material Ca3Ru2O7.
-While its ground state lacks AM features, alternative magnetic configurations enable their realization.
-AM arises in spin arrangements connected through rotational and translational symmetries.
-In particular, configuration B, which alternates spin orientation both in-plane and between layers, exhibits spin connectivity via rotations along the -axis and translations in the - plane.
-This configuration hosts a P-2 -wave altermagnetic state, which is orbital-selective, with AM primarily arising from the - orbitals.
-When the spin-orbit coupling is included, with the Néel vector oriented along the -axis, time-reversal symmetry breaking leads to weak ferromagnetism along the -axis due to the staggered Dzyaloshinskii-Moriya interaction [7, 15].
-While configuration B has not been experimentally observed in pristine Ca3Ru2O7, it has been reported in doped systems, such as Ca3(TixRu1-x)2O7, suggesting that this system could host an AM ground state under suitable conditions.
-Additionally, our results indicate that strain plays a crucial role in tuning AM properties.
-Tensile strain along the -axis enhances AM band splitting by approximately (for strain), whereas compressive strain suppresses it by .
-Moreover, we predict a strain-driven phase transition from an AFM to an AM state (configuration A to B) under compressive strain exceeding .
-While previous experiments have manipulated magnetic anisotropy using uniaxial strain (), the impact of biaxial strain remains largely unexplored.
-Strain-tuned AM phases could significantly expand the known quantum phenomena in the Can+1RunO3n+1 family, which already exhibit strain-driven transitions, such as Mott phase suppression in Ca2RuO4 [17], the emergence of a Kondo effect, and a metal-to-semiconductor transition in CaRuO3 ( strain) [42].
-Our findings provide a fundamental step toward uncovering possible AM states in the Ca3Ru2O7 family and pave the way for future theoretical and experimental investigations of biaxial strain as a tool for engineering AM phases in correlated oxides.
+S7 illustrate how freezing the atomic positions primarily alters the gap structure rather than the AM spin splitting itself, reinforcing the observation that the observed strain-dependent AM behavior is primarily dictated by electronic exchange interactions, with octahedral distortions acting as a secondary fine-tuning mechanism.
+Furthermore, Fig.
+8(d) reveals how the RuO octahedral distortion angle responds to biaxial strain in each magnetic configuration.
+In configuration A, increases smoothly from to strain.
+Configuration B shows a more intricate pattern: under compressive strain (), increases as the lattice compresses in the plane.
+This response reflects stronger electronic localization in the antiferromagnetic arrangement (see Fig.
+3(b)), which preserves magnetic exchange by enhancing octahedral distortion.
+Under tensile strain, in-plane expansion tends to relax distortions while the slight contraction along the axis reinforces them.
+The net change in depends on the balance between these two opposing effects.
+The evolution of in configuration B correlates directly with the energy profile in Fig.
+8(a).
+Above compressive strain, configuration B becomes lower in energy than configuration A and increases sharply.
+This coincidence demonstrates that enhanced octahedral distortion drives the AFM-to-AM transition.
+Although such strain levels are challenging to achieve in bulk crystals, epitaxial oxide films can routinely accommodate comparable lattice mismatches, providing a practical platform to realize and probe distortion-driven magnetic and electronic phase transitions in layered perovskites [45, 46, 47, 48].
+This work investigates the emergence of altermagnetic states in the correlated oxide Ca3Ru2O7.
+While its ground state is not altermagnetic, symmetry-allowed AM order arises in alternative spin configurations.
+In the non-relativistic case, Configuration B, characterized by a Néel-type spin pattern, hosts a P-2 -wave altermagnetic state.
+Including spin-orbit coupling with the Néel vector along induces additional symmetry breaking, allowing a -wave component in the spin texture.
+This hybridization between - and -wave features may result from the absence of inversion symmetry and the presence of nonsymmorphic operations.
+A complete understanding of the -wave origin will require further analysis based on spin space group symmetries under non-collinear magnetism.
+Although configuration B has not been reported in pristine Ca3Ru2O7, it appears in doped compounds such as Ca3(TixRu1-x)2O7, suggesting that suitable conditions can stabilize the AM state.
+Our results further demonstrate that biaxial strain acts as an effective tuning parameter for the altermagnetic band splitting: tensile strain (or equivalently, compression along the -axis) enhances the splitting by approximately at strain, whereas compressive strain reduces it by about .
+We also predict a strain-induced transition from the AFM ground state to the AM state beyond 2% compressive strain.
+While prior studies have used uniaxial strain to manipulate magnetic anisotropy [26], the effects of biaxial strain remain largely unexplored.
+Inducing altermagnetism in Ca3Ru2O7 would create a rare noncentrosymmetric altermagnetic material and provide a platform to explore the interplay between odd- and even-wave spin-momentum locking.
+Identifying such phases would broaden the spectrum of quantum phenomena known in the Can+1RunO3n+1 family, which already exhibits strain-tunable transitions, including Mott suppression in Ca2RuO4 ( strain)[18], the emergence of a Kondo effect, and a metal-to-semiconductor transition in CaRuO3 ( strain)[45].
+Our results highlight biaxial strain as a promising route to control altermagnetic phases in Can+1RunO3n+1-related materials, aligning with recent proposals to manipulate antiferromagnetic-altermagnetic phase transitions via strain [49, 50].
+These findings motivate further exploration of symmetry-driven spin textures in systems with strong spin-orbit coupling.
+Note: Recent studies have shown that SOC can induce interesting phenomena such as polarization switching in Ca3Mn2O7, driven by magnetic anisotropy [43].
+Given its structural similarity to Ca3Ru2O7, our results suggest that Ca3Mn2O7 may also host mixed-symmetry altermagnetic order under SOC.
+Further exploration of this system could offer new pathways for tuning altermagnetic behavior via SOC.
+Methods
+We perform a theoretical analysis using the density functional theory with and without spin-orbit coupling.
+We use the plane-wave pseudopotential method implemented in the Vienna ab-initio simulation package (VASP)[51] within the generalized gradient approximations (GGA), and it gives structural properties closer to the experimental values.
+The electronic valence considered are: Ru: 54 and O: 22, and Ca: 334.
+We use a plane-wave energy cutoff of 650 eV and set a Regular Monkhorst-Pack grid of 775 to perform the atomic relaxation and 11117 to perform the self-consistent calculation.
+We use a fine k-grid of 14147 within the tetrahedron method for the density of states.
+We perform the structural optimization of the unit cell until a force convergence threshold of at least 10-3 eV/Å per atom.
+To account for electronic correlations in the Ru -orbitals, we use the Dudarev approach [52] with an effective Hubbard parameter eV, previously shown to yield accurate structural and electronic properties for CRO [36].
+We use a Cartesian coordinate system with crystallographic axes aligned as , , and .
+Data availability
+All data supporting the findings of this study are included in the article and its Supplementary Information.
+Code Availability
+Density functional theory calculations were performed using the Vienna ab initio Simulation Package (VASP), accessed under an institutional license through the high-performance computer at the NHR Center of TU Dresden and on the high-performance computers Noctua 2 at the NHR Center PC2.
+Custom Python scripts developed for post-processing and analysis are available from the corresponding author upon reasonable request.
 Acknowledgments
@@
 is supported by the Foundation for Polish Science project “MagTop” no.
-FENG.02.01-IP.05-0028/23 co-financed by the European Union from the funds of Priority 2 of the European Funds for a Smart Economy Program 2021–2027 (FENG).
+FENG.02.01-IP.05-0028/23 co-financed by the European Union from the funds of Priority 2 of the European Funds for a Smart Economy Program 2021-2027 (FENG).
 J.W.G.
@@
 HE 3543/42-1.
+The authors gratefully acknowledge the computing time made available to them on the high-performance computer at the NHR Center of TU Dresden and on the high-performance computers Noctua 2 at the NHR Center PC2.
+Author Contributions
+A.L.
+and J.G.
+performed the calculations and contributed to the organization of the article.
+A.L.
+and C.A.
+held foundational discussions about the study.
+A.L., J.G., and C.A.
+carried out the symmetry analysis.
+A.L.
+and T.B.
+performed the electronic structure analysis.
+A.L., J.G., and C.A.
+wrote the initial draft.
+All authors contributed to the discussion of the results and collaborated in editing and shaping the final version of the manuscript.
+Competing interests
+The authors declare no competing interests.

### v3

@@
 Although these magnetic states have not been experimentally observed in CRO, similar phases have been reported in doped Ca3Ru2O7 systems, such as Ti-doped Ca3(Ru1-xTix)2O7 (configuration B) [33], and in the sister compound Sr3Ru2O7 (configuration D) [34].
-Notably, variations in magnetic spin arrangements lead to significant structural and electronic changes (further structural details are given in supplementary information file (SI) table S1).
+Notably, variations in magnetic spin arrangements lead to significant structural and electronic changes (further structural details are given in SM table S1).
 Specifically, FM and AFM in-plane arrangements result in metallic (A, C) and a narrow insular state (B, D), respectively.
 Next, we will focus on the AM phases of CRO, primarily in Configuration B.
-In addition to the electronic differences, configuration B exhibits a lower magnetic moment than configuration A (details on Table S1).
+| E | E-State | M-Phase | () | |
+| A | 0 | Metallic | AFM | 1.44 |
+| B | 34.90 | N-Insulator | AM | 1.35 |
+| C | 36.46 | Metallic | AM | 1.38/1.45 |
+| D | 39.22 | N-Insulator | AFM | 1.35 |
+In addition to the electronic differences, configuration B exhibits a lower magnetic moment than configuration A (details on Table 1).
 In configuration B, the magnetic moments form a Néel-type collinear ordering with alternating spin orientations.
@@
 The resulting spectral redistribution lowers the chemical potential relative to the less-affected, half-filled and orbitals.
-The gap thus reflects an orbital-selective mechanism driven by the interplay between magnetism and crystal-field effects, as proposed for Ca2RuO4 [40]
+The gap thus reflects an orbital-selective mechanism driven by the interplay between magnetism and crystal-field effects, as proposed for Ca2RuO4 [39].
 Figure 4 shows the non-relativistic band structure for the A and B magnetic configurations.
@@
 In contrast, the bands between eV, dominated by the orbital character, show negligible spin splitting (see, Fig.
-S3).
+LABEL:fig:S2).
 Therefore, configuration B exhibits orbital-selective altermagnetism with the altermagnetism present in the - bands but not in the bands, this is similar to that observed in Ca2RuO4 [39].
@@
 C displays different magnetic states in which all the t2g orbitals contribute to the AM character; therefore, it does not exhibit orbital selective AM, see more details in Fig.
-S3.
+LABEL:fig:S2.
 The lack of inversion symmetry in Ca3Ru2O7 enables sizable SOC effects, which play a central role in its magnetic properties [27].
@@
 Fig.
-S6 shows the evolution of the spin splitting across the full -path and occupied bands under different strain conditions (see Brillouin zone in Fig.
+S9 shows the evolution of the spin splitting across the full -path and occupied bands under different strain conditions (see Brillouin zone in Fig.
 4).
@@
 All computational details and equations are provided in the Supplementary Information (Fig.
-S7).
+S10).
+To further dissect the symmetry of the spin splitting in three dimensions (details provided in the SM, section SVI), we perform a multipolar analysis of the energy texture .
+In the non-relativistic limit, the response is purely even in , confirming the expected pure -wave altermagnetic state dictated by symmetry.
+Including SOC activates odd-parity components and transforms the system into a hybrid state with concurrent -, -, and -wave character.
+Using spinor band pairs near the Fermi level, we find overall -wave dominance, with fractional -norm weights , , and .
+The hybridization is strongly anisotropic across spin channels: the and components are -dominated (, ), whereas the component is -dominated ().
 Discussion
@@
 Conversely, under tensile strain, the lattice is stretched in the -plane while slightly contracting along , effectively narrowing the bandwidth and reinforcing the altermagnetic character (Fig.
-S5).
+S8 SM).
 This effect is reflected in the rising AMQ values in Fig.
@@
 In particular, the density of states in Fig.
-S5 confirms that compressive strain broader states, reducing AM features.
+S8 confirms that compressive strain broader states, reducing AM features.
 In contrast, tensile strain narrows the bandwidth, reinforcing the AM character.
 Moreover, the band structures shown in Fig.
-S7 illustrate how freezing the atomic positions primarily alters the gap structure rather than the AM spin splitting itself, reinforcing the observation that the observed strain-dependent AM behavior is primarily dictated by electronic exchange interactions, with octahedral distortions acting as a secondary fine-tuning mechanism.
+S10 illustrate how freezing the atomic positions primarily alters the gap structure rather than the AM spin splitting itself, reinforcing the observation that the observed strain-dependent AM behavior is primarily dictated by electronic exchange interactions, with octahedral distortions acting as a secondary fine-tuning mechanism.
 Furthermore, Fig.
@@
 We use a Cartesian coordinate system with crystallographic axes aligned as , , and .
-Data availability
-All data supporting the findings of this study are included in the article and its Supplementary Information.
-Code Availability
-Density functional theory calculations were performed using the Vienna ab initio Simulation Package (VASP), accessed under an institutional license through the high-performance computer at the NHR Center of TU Dresden and on the high-performance computers Noctua 2 at the NHR Center PC2.
-Custom Python scripts developed for post-processing and analysis are available from the corresponding author upon reasonable request.
 Acknowledgments
@@
 HE 3543/42-1.
-The authors gratefully acknowledge the computing time made available to them on the high-performance computer at the NHR Center of TU Dresden and on the high-performance computers Noctua 2 at the NHR Center PC2.
 Author Contributions
@@
 All authors contributed to the discussion of the results and collaborated in editing and shaping the final version of the manuscript.
-Competing interests
+1 Competing interests
 The authors declare no competing interests.
+Supplementary Information
+Hybrid d/p-wave altermagnetism in Ca3Ru2O7 and strain-controlled spin splitting
+SI Ca3Ru2O7 structural parameters
+Table S1 shows the energy difference regarding the ground state system (Configuration A) considering spin orbit coupling (SOC) and without SOC (NSOC), respectively (considering the spins along direction).
+, and are the lattice parameter (obtained from NSOC calculations).
+| Configuration | ENSOC (meV/Ru) | ESOC (meV/Ru) | a/b (Å) | c (Å) |
+| A | 0 | 0 | 5.415/5.601 | 19.54 |
+| B | 34.90 | 36.86 | 5.415/5.726 | 19.23 |
+| C | 36.46 | 60.55 | 5.398/5.652 | 19.449 |
+| D | 39.22 | 37.32 | 5.419/5.722 | 19.261 |
+| Exp [1] | – | – | 5.367/5.535 | 19.521 |
+Note: The calculations including spin-orbit coupling were performed using the relaxed structures obtained from collinear calculations and do not include further structural relaxation.
+Effect Hubbard-U energy differences: Table S2 shows the energy difference relative to the ground state (Configuration A) for three different values of the Hubbard parameter.
+We observe that the choice of does not significantly affect the relative stability of the magnetic phases with respect to configuration A.
+For eV, configuration A begins to exhibit structural changes, with distortions in the octahedral environment indicating the onset of structural instability.
+This behavior has been previously investigated [2], and it has been established that at eV the system stabilizes in a different structural phase [3, 2].
+| Configuration | E (U = 0.5) | E (U = 1.0) | E (U = 1.1) |
+| A | 0 | 0 | 0 |
+| B | 48.41 | 34.90 | 30.76 |
+| C | 52.35 | 36.46 | 32.79 |
+| D | 74.0 | 39.23 | 37.29 |
+SII Density of states with and without Hubbard-U interaction.
+Figure S1 shows the total density of states (DOS) for configurations A and B, with and without the inclusion of the Hubbard-U interaction.
+At U=0 eV, both systems exhibit a metallic state.
+When the Hubbard-
+U interaction is introduced, configuration B favors a narrow-gap insulating state, whereas configuration A remains metallic.
+This different response to the Hubbard-U
+interaction arises from the localization versus delocalization effects associated with the symmetry of the wave functions, which in turn are determined by the magnetic configurations.
+Specifically, configuration B corresponds to a singlet state, where the spin part of the wave function is antisymmetric under exchange, leading to a symmetric (bonding-like) spatial part that favors electron delocalization.
+In contrast, configuration A resembles a triplet state, with a symmetric spin part and thus an antisymmetric (antibonding-like) spatial part, resulting in more localized electron clouds.
+In essence, the Hubbard- term promotes electron localization.
+Since configuration A already exhibits a relatively localized electronic character, the effect of a small value is minimal.
+However, by increasing , a gap opens, leading to an insulating state with a narrow gap, similar to that observed in configuration B.
+More details about the interplay among the different degrees of freedom in configuration A can be found in Ref.
+[3].
+In contrast, configuration B initially favors a more delocalized state; therefore, the introduction of induces localization, which perturbs the spatial part of the wave function and, consequently, alters the magnetic order.
+SIII Ca3Ru2O7 altermagnetic configurations
+Fig. S2 (upper panel) shows the band structure for the AM configurations of Ca3Ru2O7: configuration B (a) and configuration C (b).
+Fig.
+LABEL:fig:S2 shows the band structure projected onto the -orbitals by occupation state for the B and C configurations.
+While the configuration B hosts a P-2 dyz-wave altermagnetism, the configuration C hosts a dxz-wave altermagnetism.
+Configuration C exhibits altermagnetism in both dxy and
+dxz-dyz bands, therefore, the altermagnetism of the configuration C is not orbital selective. In the case of the B configuration, the main contributions to the AM bands are between -1.5 and -1.0 eV.
+SIV Relativistic bands
+Figure LABEL:fig:S4 shows the band structure for configuration B.
+Upper panel: Band structure along the path, where altermagnetic splitting is most pronounced (see Fig.
+4 in the main text).
+The projection retains strong AM features at lower energies, consistent with the collinear case.
+In contrast, the projection reveals AM features between eV and eV.
+Lower panel: Band structure along the path. The and components retain the antiferromagnetic character, while the projection shows clear AM features near the Fermi level.
+SV Band structure under strain
+Figure LABEL:fig:S4 shows the variation of the DOS at the Fermi level as a function of strain.
+It can be observed that under compressive strain, the DOS becomes broader, while under tensile strain, it becomes more localized.
+Figure S8 presents the band structure under different strain values (-plane) for two cases.
+The upper panels correspond to fully relaxed structures, while the lower panels depict the same strain conditions (same lattice parameters) but with atomic positions fixed at their zero-strain values (non-relaxed calculations).
+In this scenario, the systems retain the Ru-O rotation from the zero-strain configuration.
+Notably, removing the Ru-O distortions primarily affects the band gap.
+The main text focused on the absolute energy difference between majority and minority bands.
+However, in Fig.
+S10, we instead plot the signed difference (spin-up minus spin-down) to highlight how the splitting changes sign across the Brillouin zone.
+For instance, the splitting inverts going from to and from to .
+Such sign reversals showcase the complementary nature of the -dependent spin splitting that emerges in the altermagnetic state.
+While taking the absolute value is helpful to quantify the splitting strength, plotting the raw sign clarifies how the splitting evolves—and sometimes flips—across distinct points in -space.
+Figure S10 shows the resulting along the entire -path, defined by
+where is the number of bands at each -point within the specified energy range, and is the number of occupied bands.
+These plots show that is sensitive to the applied strain.
+Under compressive strain, the overall magnitude of decreases compared to the unstrained system.
+In contrast, under tensile strain, it increases, underscoring how the electronic structure and spin splitting respond to lattice deformations.
+SVI 3D Multipolar Analysis of the Spin-Resolved Splitting
+Conventions and Brillouin-zone averaging
+Let be the matrix of direct-lattice vectors (in Å) and the reciprocal-lattice matrix (in Å-1).
+Fractional momenta are mapped to Cartesian reciprocal coordinates by .
+Brillouin-zone (BZ) averages are computed discretely using the -point integration weights supplied by the electronic-structure calculation:
+Spin-resolved splitting
+In noncollinear calculations with spin–orbit coupling (SOC), Bloch states are spinors and lack a definite spin projection.
+The spin-resolved splitting for the spin component is constructed from pairs of nearby bands via a spin-contrast projection,
+where is the band-resolved spin expectation and regularizes accidental zeros in the spin contrast (values as small as are sufficient in practice).
+For a chosen set of band pairs (e.g., the topmost occupied spinor bands), the effective splitting is the arithmetic mean
+In the collinear (no-SOC) case with quantization axis , the splitting is nonzero only for :
+where is the set of bands included in the average.
+Multipolar moments ()
+The spin-resolved energy texture is analyzed through the first three BZ moments [4]:
+By parity, and are odd under and diagnose odd-parity () components, whereas is even and diagnoses the even-parity () component. All averages are evaluated on the full 3D mesh.
+Strength measures and fractional weights
+For each spin channel , the overall strength of -, -, and -wave content is quantified by the norms
+The corresponding fractional weights are
+and summarize the dominant multipolar character in each spin channel. It is often convenient to report both per-channel fractions and overall fractions [4, 5].